OpenStudy (ttop0816):

would it be 30 degrees?? Which of the following angles can be used to evaluate sec120°? 30° 45° 60°

OpenStudy (anonymous):

sec (180-60) = sec 120 hence 60 degrees would be the answer

OpenStudy (usukidoll):

120 is in quadrant 2... since we didn't reach over 180 degrees which is half a circle... subtract 180 by 120 = 180-120 = 60

OpenStudy (usukidoll):

cos 60 is 1/2 since sec 60